OpenStudy (anonymous):

Write a polynomial function in standard form with the given zeros: -1,2/3

OpenStudy (anonymous):

y = (x+1)(3x-2)

OpenStudy (anonymous):

x+1=0 -1 -1 x = -1 3x-2=0 +2 +2 3x = 2 divide by 3 on both sides x = (2/3)

OpenStudy (anonymous):

For standard form, expand the equation to the form ax+by=c

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

y=(x+1)(3x-2) is not standard, use FOIL: (x*-2)+(x*3x)+(1*3x)+(1*-2)